# Language Models Encode Collaborative Signals in Recommendation

### v1

@@ Abstract
Recent studies empirically indicate that language models (LMs) encode rich world knowledge beyond mere semantics, attracting significant attention across various fields. However, in the recommendation domain, it remains uncertain whether LMs implicitly encode user preference information. Contrary to the prevailing understanding that LMs and traditional recommender models learn two distinct representation spaces due to a huge gap in language and behavior modeling objectives, this work rethinks such understanding and explores extracting a recommendation space directly from the language representation space. Surprisingly, our findings demonstrate that item representations, when linearly mapped from advanced LM representations, yield superior recommendation performance. This outcome suggests the homomorphism between the language representation space and an effective recommendation space, implying that collaborative signals may indeed be encoded within advanced LMs. Motivated by these findings, we propose a simple yet effective collaborative filtering (CF) model named AlphaRec, which utilizes language representations of item textual metadata (e.g., titles) instead of traditional ID-based embeddings. Specifically, AlphaRec is comprised of three main components: a multilayer perceptron (MLP), graph convolution, and contrastive learning (CL) loss function, making it extremely easy to implement and train. Our empirical results show that AlphaRec outperforms leading ID-based CF models on multiple datasets, marking the first instance of such a recommender with text embeddings achieving this level of performance. Moreover, AlphaRec introduces a new language-representation-based CF paradigm with several desirable advantages: being easy to implement, lightweight, rapid convergence, superior zero-shot recommendation abilities in new domains, and being aware of user intention. Codes are available at https://github.com/LehengTHU/AlphaRec.
@@ 1 Introduction
(a) Linearly mapping language representations into the recommendation space
@@
Figure 1: Linearly mapping item titles in language representation space into recommendation space yields superior recommendation performance on Movies & TV [1] dataset. (1(a)) The framework of linear mapping. (1(b)) The recommendation performance comparison between leading CF recommenders and linear mapping. (1(c)) The t-SNE [2] visualizations of movie representations, with colored lines linking identical movies or user intention across language space (left) and linearly projected recommendation space (right).
@@
Language models (LMs) have achieved great success across various domains [3, 4, 5, 6, 7], prompting a critical question about the knowledge encoded within their representation spaces. Recent studies empirically find that LMs extend beyond semantic understanding to encode comprehensive world knowledge about various domains, including game states [8], lexical attributes [9], and even concepts of space and time [10] through language modeling. However, in the domain of recommendation where the integration of LMs is attracting widespread interest [11, 12, 13, 14, 15], it remains unclear whether LMs inherently encode relevant information on user preferences and behaviors. One possible reason is the significant difference between the objectives of language modeling for LMs and user behavior modeling for recommenders [16, 17, 18, 19].
@@
Currently, one prevailing understanding holds that general LMs and traditional recommenders encode two distinct representation spaces: the language space and the recommendation space (i.e., user and item representation space), each offering potential enhancements to the other for recommendation tasks [17, 20]. On the one hand, when using LMs as recommenders, aligning the language space with the recommendation space could significantly improve the performance of LM-based recommendation [14, 21, 22, 23]. Various alignment strategies are proposed, including fine-tuning LMs with recommendation data [15, 16, 24, 25, 26], incorporating embeddings from traditional recommenders as a new modality of LMs [17, 20, 27], and extending the vocabulary of LMs with item tokens [18, 19, 28, 29, 30, 31]. On the other hand, when using LMs as the enhancer, traditional recommenders greatly benefit from from leveraging text representations [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45], semantic and reasoning information [46, 47, 48, 49], and generated user behaviors [50, 51]. Despite these efforts, explicit explorations of the relationship between language and recommendation spaces remain largely unexplored.
@@
In this work, we rethink the prevailing understanding and explore whether LMs inherently encode user preferences through language modeling. Specifically, we test the possibility of directly deriving a recommendation space from the language representation space, assessing whether the representations of item textual metadata (e.g., titles) obtained from LMs can independently achieve satisfactory recommendation performance. Positive results would imply that user behavioral patterns, such as collaborative signals (i.e., user preference similarities between items) [52, 53], may be implicitly encoded by LMs. To test this hypothesis, we employ linear mapping to project the language representations of item titles into a recommendation space (see Figure 1(a)). Our observations include:
@@
  • Surprisingly, this simple linear mapping yields high-quality item representations, which achieve exceptional recommendation performance (see Figure 1(b) and experimental results in Section 2).
  • The clustering of items is generally preserved from the language space to the recommendation space (see Figure 1(c)). For example, movies with the theme of superheroes and monsters are gathering in both language and recommendation spaces.
@@
  • Interestingly, the linear mapping effectively reveals preference similarities that may be implicit or even obscure in the language space. For instance, while certain movies, such as those of homosexual movies (illustrated in Figure 1(c)), show dispersed representations in the language space, their projections through linear mapping tend to cluster together, reflecting their genres affiliation.
These findings indicate the homomorphism [54] between the language representation space of LMs and an effective item representation space for recommendation. Motivated by this insight, we propose a new language-representation-based recommendation paradigm for general collaborative filtering (CF), which utilizes the pre-trained language representations of item titles as the item input and the average historical interactions’ representations as the user input. Different from traditional ID-based CF models [55, 56, 52] that heavily rely on trainable user and item IDs, this paradigm solely uses pre-trained LM embeddings and completely abandons ID-based embeddings. In this paper, to fully explore the potential of advanced language representations, we adopt a simple model architecture consisting of a two-layer MLP with graph convolution, and the popular contrastive loss, InfoNCE [57, 58, 59], as the objective function. This model is named AlphaRec for its originality, a series of good properties, and an adjustable hyperparameter $\alpha$ for user intention capture.
@@
Benefiting from paradigm shifts from ID-based embeddings to language representations, AlphaRec presents three desirable advantages. First, AlphaRec is notable for its simplicity, lightweight, rapid convergence, and exceptional recommendation performance (see Section 4.1). We empirically demonstrate that, for the first time, such a simple model with embeddings from pre-trained LMs can outperform leading CF models on multiple datasets. This finding strongly supports the possibility for developing language-representation-based recommender systems. Second, AlphaRec exhibits a strong zero-shot recommendation capability across untrained datasets (see Section 4.2). By co-training on three Amazon datasets (Books, Movies & TV, and Video Games) [1], AlphaRec can achieve performance comparable to the fully-trained LightGCN on entirely different platforms (MovieLens-1M [60] and BookCrossing [61]), and even exceed LightGCN in a completely new domain (Amazon Industrial), without additional training on these target datasets. This capability underscores AlphaRec’s potential to develop more general recommenders. Third, AlphaRec is user-friendly, offering a new research paradigm that enhances recommendation by leveraging language-based user feedback (see Section 4.3). Endowed with its inherent semantic comprehension of language representations, AlphaRec can refine recommendations based on user intentions expressed in natural language, enabling traditional CF recommenders to evolve into intention-aware systems through a straightforward paradigm shift.
@@ 2 Uncovering Collaborative Signals in LMs via Linear Mapping
In this section, we aim to explore whether LMs implicitly encode user preference similarities in their representation spaces. We first formulate the personalized item recommendation task, then detail the linear mapping method and its empirical findings. Empirical evidence indicates the homomorphism [54] between the representation spaces of advanced LMs and effective recommendation spaces.
Task formulation. Personalized item recommendation with implicit feedback aims to select items $i\in\mathcal{I}$ that best match user $u$’s preferences based on binary interaction data $\mathbf{Y}=[y_{ui}]$, where $y_{ui}=1$ ($y_{ui}=0$) indicates user $u\in\mathcal{U}$ has (has not) interacted with item $i$ [59]. The primary objective of recommendation is to model the user-item interaction matrix $\mathbf{Y}$ using a scoring function $\hat{y}:\mathcal{U}\times\mathcal{I}\rightarrow\mathbb{R}$, where $\hat{y}_{ui}$ measures $u$’s preference for $i$. The scoring function $\hat{y}_{ui}=s\circ\phi_{\theta}(\mathbf{x}_{u},\mathbf{x}_{i})$ comprises three key components: pre-existing features $\mathbf{x}_{u}$ and $\mathbf{x}_{i}$ for user $u$ and item $i$, a representation learning module $\phi_{\theta}(\cdot,\cdot)$ parametrized by $\theta$, and a similarity function $s(\cdot,\cdot)$. The representation learning module $\phi_{\theta}$ transfers $u$ and $i$ into representations $\mathbf{e}_{u}$ and $\mathbf{e}_{i}$ for similarity matching $s(\cdot,\cdot)$, and the Top-$K$ highest scoring items are recommended to $u$.
Different recommenders employ various pre-existing features $\mathbf{x}_{u},\mathbf{x}_{i}$ and representation learning architecture $\phi_{\theta}(\cdot,\cdot)$. Traditional ID-based recommenders use one-hot vectors as pre-existing features $\mathbf{x}_{u},\mathbf{x}_{i}$. The choice of ID-based representation learning architecture $\phi_{\theta}$ can vary widely, including ID-based embedding matrix [55], multilayer perception [62], graph neural network [52, 63], and variational autoencoder [64]. We adopt a commonly used similarity function, cosine similarity [65, 58] $s(\mathbf{e}_{u},\mathbf{e}_{i})=\frac{{\mathbf{e}_{u}}^{\top}\mathbf{e}_{i}}{\left\lVert\mathbf{e}_{u}\right\rVert\cdot\left\lVert\mathbf{e}_{i}\right\rVert}$, in this paper.
@@
Linear mapping. Linear mapping ensures the homomorphism [54] between the vector spaces, and is therefore widely used in the study on the representation space properties of LMs [10, 66, 67]. In this section, we adopt a trainable linear mapping matrix $\bm{W}$ as the representation learning module $\phi_{\theta}$, mapping representations from the language space into an item recommendation space. High performance on the test set would indict the homomorphism between the language space and an effective recommendation space, which further suggests the possible existence of collaborative signals (i.e., user preference similarities between items) [56] in the language representation space [68, 10]. The overall framework of the linear mapping process is illustrated in Figure 1(a). Specifically, we utilize LMs as feature extractors, leveraging the language representations of item titles as initial item feature $\mathbf{x}_{i}$. For initial user feature $\mathbf{x}_{u}$, we use the average of the title representations of historically interacted items, defined as $\mathbf{x}_{u}=\frac{1}{|\mathcal{N}_{u}|}\sum_{i\in\mathcal{N}_{u}}\mathbf{x}_{i}$, where $\mathcal{N}_{u}$ is the set of items user $u$ has interacted with. Detailed procedures for obtaining these language-based features are provided in Appendix B.2. The linear mapping matrix sets user and item representations as $\mathbf{e}_{u}=\bm{W}\mathbf{x}_{u}$ and $\mathbf{e}_{i}=\bm{W}\mathbf{x}_{i}$ respectively. To learn the linear mapping matrix $\bm{W}$, we adopt the InfoNCE loss [57] as the objective function, which has demonstrated state-of-the-art performance in both ID-based [69, 70] and LM-enhanced CF recommenders [47] (refer to Equation (4) for the formula). We directly use linearly mapped representations $\mathbf{e}_{u}$ and $\mathbf{e}_{i}$ to calculate the user-item similarity $s(\mathbf{e}_{u},\mathbf{e}_{i})$ for recommendation.
Empirical findings. We compare the recommendation performance of the linear mapping method with three classical CF baselines, matrix factorization (MF) [55, 71], MultVAE [64], and LightGCN [56] (see more details about baselines in Appendix C.2.1). We report three widely used metrics Hit Ratio (HR@$K$), Recall@$K$, Normalized Discounted Cumulative Gain (NDCG@$K$)) to evaluate the effectiveness of linear mapping, with $K$ set by default at 20. We evaluate a wide range of LMs, including BERT-style models [4, 5], decoder-only large language models (LLMs) [6, 72], and LLM-based text embedding models [73, 74] (see Appendix B.1 for details about used LMs).

### v2

# Do Language Models Encode Collaborative Signals in Recommendation? Findings and Potentials

@@ Abstract
Recent studies empirically indicate that language models (LMs) encode rich world knowledge beyond mere semantics, attracting significant attention across various fields. However, in the recommendation domain, it remains uncertain whether LMs implicitly encode user preference information. Contrary to prevailing understanding that LMs and traditional recommenders learn two distinct representation spaces due to the huge gap in language and behavior modeling objectives, this work re-examines such understanding and explores extracting a recommendation space directly from the language representation space. Surprisingly, our findings demonstrate that item representations, when linearly mapped from advanced LM representations, yield superior recommendation performance. This outcome suggests the possible homomorphism between the advanced language representation space and an effective item representation space for recommendation, implying that collaborative signals may be implicitly encoded within LMs. Motivated by the finding of homomorphism, we explore the possibility of designing advanced collaborative filtering (CF) models purely based on language representations without ID-based embeddings. To be specific, we incorporate several crucial components (i.e., a multilayer perceptron (MLP), graph convolution, and contrastive learning (CL) loss function) to build a simple yet effective model, with the language representations of item textual metadata (i.e., title) as the input. Empirical results show that such a simple model can outperform leading ID-based CF models on multiple datasets, which sheds light on using language representations for better recommendation. Moreover, we systematically analyze this simple model and find several key features for using advanced language representations: a good initialization for item representations, superior zero-shot recommendation abilities in new datasets, and being aware of user intention. Our findings highlight the connection between language modeling and behavior modeling, which can inspire both natural language processing and recommender system communities  [^1] Codes are available at https://github.com/LehengTHU/AlphaRec.
@@ 1 Introduction
(a) Linearly mapping language representations into the behavior space for recommendation
@@
Figure 1: Linearly mapping item titles in language representation space into behavior space yields superior recommendation performance on Movies & TV (Ni et al., 2019) dataset. (1(a)) The framework of linear mapping. (1(b)) The recommendation performance comparison between leading CF recommenders and linear mapping. (1(c)) The t-SNE (Van der Maaten & Hinton, 2008) visualizations of movie representations, with colored lines linking identical movies or user intention across language space (left) and linearly projected behavior space for recommendation (right).
@@
Language models (LMs) have achieved great success across various domains (Vaswani et al., 2017; Devlin et al., 2019; Dubey et al., 2024; OpenAI, 2023), raising a critical question about the knowledge encoded within the language space. Recent studies empirically find that LMs extend beyond semantic understanding to encode comprehensive world knowledge about various domains, such as game states (Li et al., 2023a), lexical attributes (Vulic et al., 2020), and even concepts of space and time (Gurnee & Tegmark, 2023) through language modeling. However, in the domain of recommendation where the integration of LMs is attracting widespread interest (Fan et al., 2023; Li et al., 2023b; Wu et al., 2023a), it remains unclear whether LMs inherently encode relevant information on user preferences and behaviors in the language space.
@@
Currently, one prevailing understanding holds that general LMs and traditional recommenders (e.g., collaborative filtering models (Koren et al., 2009; He et al., 2021)) encode distinct representation spaces — one for language space and the other for behavior space — but they offer the potential to enhance each other in downstream recommendation tasks (Liao et al., 2024). Specifically, on the one hand, when using LMs as recommenders to directly output items of interest, aligning the language space with the behavior space can significantly improve the recommendation performance (Lin et al., 2023a; Vats et al., 2024; Xu et al., 2024). Various alignment strategies are proposed, including fine-tuning LMs with user behavior data (Zhang et al., 2023d; Bao et al., 2023; Geng et al., 2022; Cui et al., 2022; Lin et al., 2023b), incorporating embeddings from traditional recommenders as a new modality of LMs (Liao et al., 2024; Zhang et al., 2023e; Yang et al., 2023), and extending the vocabulary of LMs with item tokens (Zhu et al., 2023; Zheng et al., 2023; Rajput et al., 2023; Zhai et al., 2024). On the other hand, when using LMs as the enhancer to represent item content (e.g., text metadata), traditional recommenders greatly benefit from text embeddings (Yuan et al., 2022; 2023; Li et al., 2023c; Hou et al., 2024a; Liu et al., 2024a), semantic and reasoning information (Wei et al., 2024; Ren et al., 2024b; Xi et al., 2023), and generated user behaviors (Zhang et al., 2023b; c). Despite these efforts, explicit studies of the relationship between language and behavior spaces remain largely unexplored in the recommendation domain.
@@
In this work, we re-examine this prevailing understanding, by exploring whether LM-generated language space has inherently encoded user preferences and behaviors. Specifically, we test the possibility of directly deriving a behavior space from the language space — that is, we assess whether the language representations of item text metadata (e.g., titles) generated by LMs can independently predict user behaviors and achieve competitive recommendation performance. Positive results would imply that user behavioral patterns, such as collaborative signals (i.e., users’ preference on items reflected by their behavioral similarities) (Wang et al., 2019b), might be implicitly encoded by LMs. To test this hypothesis, we employ linear mapping (Merullo et al., 2023) to project language representations of item titles into a behavior space for recommendation, as Figure 1(a) shows. Our empirical observations and findings include:
  • Before linear mapping, language representation similarities (i.e., semantic textual similarities (STS) (Muennighoff et al., 2023)) may reflect user preference similarities for item contents. Considering Figure 1(c) as an example, movies with themes of superheroes and monsters cluster together in both language and behavior spaces.
@@
  • After linear mapping, language representations are transformed into high-quality behavior representations, which achieve exceptional recommendation performance, as Figure 1(b) and experimental results in Section 3.2 show. Moreover, the performance improves as the language model size increases and remains relatively robust to prompt disturbances (see Section 3.2).
@@
  • Post-mapping language representations encode user behaviroal similarities beyond STS. For instance, while certain movies, such as those of homosexual movies (illustrated in Figure 1(c)), show low STS and their representations disperse in the language space, their projections through linear mapping tend to cluster together, reflecting high user preference similarities.
These findings suggest the homomorphism (Dieudonne, 1969) between the LM-generated language space and an expressive behavior space for recommendation. Motivated by this insight, we explore the possibility of building advanced collaborative filtering (CF) models based solely on the language space. To be specific, considering text metadata solely as items’ pre-existing features rather than the widely-used ID information, we perform a frozen LM to create the language representations of items; consequently, we apply a trainable projector (i.e., a two-layer MLP with graph convolution) to map them into a behavior space, and then employ a contrastive loss (i.e., InfoNCE (van den Oord et al., 2018; Wu et al., 2022)) to optimize. We term this model AlphaRec for its simplicity and a series of good properties. Surprisingly, our empirical results show that such a simple model can outperform leading ID-based CF models on multiple datasets. This result sheds light on using language representations for better recommendation.
@@
Furthermore, we systematically analyze this simple model and discover several potentials of adopting language representations for recommendation. First, language representations may serve as a good initialization for item representations, with few adjustments to achieve high recommendation performance (see Section 5.1). This is evidenced by the rapid training convergence of AlphaRec. Second, advanced language representations provide strong zero-shot recommendation capability across entirely new datasets (see Section 5.2). By co-training on multiple datasets, AlphaRec can achieve performance comparable to or even surpassing the fully-trained LightGCN (He et al., 2021) on new datasets without additional training. This underscores the potential of adopting advanced language representations to develop more generalizable recommenders. Third, advanced language representations provide opportunities for perceiving user intentions to refine recommendation results (see Section 5.3). Endowed with the inherent semantic comprehension of language representations, AlphaRec can adjust recommendations according to text-based user intentions, enabling recommenders to evolve into intention-aware systems through a straightforward paradigm shift.
2 Preliminary
2.1 Task Formulation
Personalized recommendation aims to learn user preferences from historical behaviors (i.e., historical interactions with items like view, click, purchase) and find items of interest to trigger users’ future behaviors (Zhou et al., 2018). In this paper, we consider one common recommendation setting: collaborative filtering (CF) (Koren et al., 2022). It aims to select item $i\in\mathcal{I}$ that best matches user $u$’s preferences based on binary interaction behaviors $\mathbf{Y}=[y_{ui}]$, where $y_{ui}=1$ indicates user $u\in\mathcal{U}$ has interacted with item $i$, and $y_{ui}=0$ otherwise (Rendle, 2022). Scrutinizing leading CF models, we summarize a common paradigm $\hat{y}_{ui}=s\circ\phi_{\theta}(\mathbf{x}_{u},\mathbf{x}_{i})$ involving three components:
  • For a user-item pair $(u,i)$, we first get their pre-existing features $\mathbf{x}_{u}$ and $\mathbf{x}_{i}$, which are usually set as ID information or one-hot encodings in CF (Koren et al., 2009; Rendle, 2022; He et al., 2021).
  • Upon $\mathbf{x}_{u}$ and $\mathbf{x}_{i}$, the representation generation module $\phi$ parameterized by $\theta$ is adopted to transfer them into behavior representations $\mathbf{e}_{u}$ and $\mathbf{e}_{i}$, encoding the behavioral patterns of users. Its architecture can vary widely, including ID-based embeddings (Koren et al., 2009), multilayer perceptions (He et al., 2017), graph neural networks (Wang et al., 2019b; Cai et al., 2023), and variational autoencoders (Liang et al., 2018).
  • Upon $\mathbf{e}_{u}$ and $\mathbf{e}_{i}$, the scoring function $s$ is used to quantify their relevance reflecting how likely user $u$ will interact with item $i$. One widely-used function is cosine similarity, $s(\mathbf{e}_{u},\mathbf{e}_{i})=\frac{{\mathbf{e}_{u}}^{\top}\mathbf{e}_{i}}{\left\lVert\mathbf{e}_{u}\right\rVert\cdot\left\lVert\mathbf{e}_{i}\right\rVert}$ (Chen et al., 2023; Wu et al., 2022).
2.2 Item Representation Generation
Here we emphasize the critical role of item representation generation, which involves transforming item $i$’s pre-existing features $\mathbf{x}_{i}$ into representations $\mathbf{e}_{i}$ suitable for recommendation. This process is essential, as the quality of these representations directly impacts the recommendation performance. In this paper, we focus mainly on two kinds of item representation generation tailor-made for different pre-existing features: ID- and LM-based generators.
ID-based generator. Prevailing CF models (Koren et al., 2009; Rendle, 2022; Koren et al., 2022; Wang et al., 2019b; He et al., 2021; Yu et al., 2024) typically convert the ID information of each item $i$ into one-hot encodings (e.g., pre-existing features $\mathbf{x}_{i}$). These sparse features are then passed through a trainable generator, such as ID embedding matrices (Koren et al., 2009; Rendle, 2022) or optionally combined with graph convolution layers (Wang et al., 2019b; He et al., 2021; Yu et al., 2024), to generate dense representations $\mathbf{e}_{i}$. Optimizing the learning of ID-based representations allows the generator to effectively learn user preferences and behaviors, leading to competitive recommendation performance. However, such ID-based generators suffer several problems, such as poor domain transferability and lack of user intention-aware abilities, since one-hot encodings lack sufficient semantics beyond being identifiers (He et al., 2021).
LM-based generator. Beyond ID information, another research line (Pazzani & Billsus, 2007; Covington et al., 2016; Liu et al., 2024b; Zhang et al., 2024a; Liu et al., 2023b) explores using the text metadata of item $i$ (e.g., titles, descriptions) as pre-existing features $\mathbf{x}_{i}$. These features are fed into the LM-based generator, typically a combination of two subsequent components: (1) A frozen LM to extract $i$’s language representation $\mathbf{z}_{i}$, such as the encoder-only LMs like BERT-style models (Devlin et al., 2019; Liu et al., 2019), the decoder-only LMs like Llama-style autoregressive models (Touvron et al., 2023b; Jiang et al., 2023) and OpenAI text embedding models (Neelakantan et al., 2022); and (2) A trainable projector to map $\mathbf{z}_{i}$ into the final representation $\mathbf{e}_{i}$, often using layers like graph convolution layers (He et al., 2021). Although such LM-based generators have been explored to enrich the item representations in literatures (Yuan et al., 2023; Li et al., 2023c; Ren et al., 2024b), few studies have demonstrated that they can solely outperform ID-based generators in recommendation tasks. Worse still, the relationship between the LM-based language space and the behavior space remain largely unexplored in the recommendation domain.
@@ 3 Uncovering Collaborative Signals in LMs via Linear Mapping
In this section, we first explore the following research questions. RQ1: Do LMs inherently encode collaborative signals (i.e., users’ preferences for items as reflected by behavioral similarities) within their representation spaces? RQ2: If so, does the presence of such signals scale with model size, and are they robust across different settings? To investigate these questions, we use linear mapping to project language representations of item titles into a behavior space for recommendation. We detail the implementation of the linear mapping in Section 3.1. Subsequently, in Section 3.2, we empirically assess the existence and robustness of collaborative signals in language representations.
3.1 Linear Mapping
Linear mapping is effective to study the representation properties of LMs (Merullo et al., 2023; Alain & Bengio, 2017), discovering the homomorphism (Dieudonne, 1969) between the language space and another space in the target domain. However, its application in the recommendation domain remains largely underexplored.
@@ 3.1 Linear Mapping
To bridge this gap, we train a linear mapping matrix $\bm{W}$ to project representations from the language space into a behavior space for recommendation. High performance of this linear mapping on the test set would indicate the presence of homomorphism between the language space and an effective behavior space, suggesting the possible existence of collaborative signals in the language representation space (Ravichander et al., 2021; Gurnee & Tegmark, 2023). The overall framework of linear mapping is illustrated in Figure 1(a). Specifically, we use frozen LMs to transform pre-existing item title features $\mathbf{x}_{i}$ into language representations $\mathbf{z}_{i}$. To derive user representations, we compute the average of the language representations of the items a user $u$ has interacted with, denoted as $\mathbf{z}_{u}=\frac{1}{|\mathcal{N}_{u}|}\sum_{i\in\mathcal{N}_{u}}\mathbf{z}_{i}$, where $\mathcal{N}_{u}$ is the set of user $u$’s historical items. See Appendix B.2 for detailed procedures for obtaining language representations. The linear mapping matrix sets behavior representations of user $u$ and item $i$ as $\mathbf{e}_{u}=\bm{W}\mathbf{z}_{u}$ and $\mathbf{e}_{i}=\bm{W}\mathbf{z}_{i}$ respectively. To optimize the matrix $\bm{W}$, we adopt the InfoNCE loss (van den Oord et al., 2018) as the objective function, which has demonstrated strong performance in both ID-based (Zhang et al., 2023a; Yu et al., 2024) and LM-based generators (Ren et al., 2024b) (refer to equation 4 for the formula).
3.2 Empirical Findings